what is the value of 12 1/8 - 9 3/5?
Answer by lenny460(1073)

what is the value of 12 1/8 - 9 3/5?

12 1/8 - 9 3/5
12 - 9 = 3
1/8 - 3/5 = 5/40 - 24/40
Take 1 from 3 and add to 1/8
Therefore;
3 - 1 = 2 or
12 - 10 = 2
1 1/8 - 3/5 = 45/40 - 24/40 = 21/40

The Answer:
2 21/40



Lennox Obuong
Algebra Student
Email: obuong3@aol.com